OpenStudy (anonymous):

Find the sum. 19x3 + (14x + 4x3) = Select one of the options below as your answer: A. 37x3 B. 37x7 C. 22x3 + 14x D. 23x3 + 14x

OpenStudy (anonymous):

19x3 + (14x + 4x3) = 19x3 + 4x3+14x =23x3 +14x

OpenStudy (mayankdevnani):

57+(14x+12)

OpenStudy (anonymous):

wat?

OpenStudy (mayankdevnani):

69+14x is answer

OpenStudy (anonymous):

Okay So On The Calculator How

OpenStudy (mayankdevnani):

i think options are wrong......

OpenStudy (anonymous):

Do I Find X

OpenStudy (mayankdevnani):

@DamnImBad

OpenStudy (anonymous):

@mayankdevnani : x is a variable and it isnt the cross symbol!!! so the ans is D

OpenStudy (mayankdevnani):

i am not do the process of cross symbol @DHASHNI

OpenStudy (mayankdevnani):

just like that 57+14x+12=69+14x

OpenStudy (anonymous):

hw did that 57 come?

OpenStudy (mayankdevnani):

oh!! 19*3=57

OpenStudy (mayankdevnani):

it is given(19*3)

OpenStudy (anonymous):

its 19 x^3

OpenStudy (anonymous):

where x is a variable

OpenStudy (anonymous):

and not a symbol

OpenStudy (anonymous):

I Already Got IT

OpenStudy (anonymous):

LOL